# Blockchain with proof of quantum work

### v2

@@ VI Experimental Implementation
Fig. 5 illustrates the mining efficiency across multiple blockchain operations. Mining efficiency is defined as the fraction of block broadcasts, excluding the genesis block, that successfully enter the strongest chain in the limit of many miners. For each curve the data is taken from 16 chains created with bootstrapping of witness distributions, each having between 512 and 1024 block broadcasts. The bootstrapping method is in agreement with results without bootstrapping reported in other figures. In Appendix D.1 we describe mining efficiency and transaction delay in greater detail, presenting additional data. The analysis compares different Chainwork definitions and post-processing methods. For the basic ±$\pm 1$ weighting method, which assigns equal weights to all block broadcasts, mining efficiency drops to close to 50% as $\mathcal{N}_{\rm zeros}\to 64$. The confidence-based weighting technique performs better, as expected, approaching 75% efficiency for $\mathcal{N}_{\rm zeros}=64$. Confidence-based Chainwork allows higher efficiency than basic +/-1 Chainwork, also with accounting of small differences in the mining rate as outlined in Appendix D.2. To calculate confidence values we used a constant $\delta W_{\alpha}$ extracted from experimental data, as detailed in Appendix D.
@@ VII Conclusions
We have proposed and successfully tested a blockchain architecture in which proof of quantum work (PoQ) is performed on quantum computers. Our approach introduces multiple quantum-based hash generation methods, with varying levels of complexity depending on the capabilities of the quantum hardware. To address the inherent probabilistic nature of quantum hash generation and validation—an unavoidable property of any quantum hash algorithm we developed techniques to ensure blockchain stability. We experimentally tested our proposed blockchain on four internationally distributed D-Wave annealing QPUs, demonstrating that despite architectural and parametric variations, these QPUs can be programmed to generate and validate each other’s hashes in a way that allowed stable blockchains to persist over thousands of broadcast blocks and hundreds of thousands of hash validations.
@@
For this demonstration, we used the critical dynamics of 3D and biclique spin glass at a fixed size as the computational problem for PoQ. This was motivated by its role in the first experimental demonstration of quantum supremacy in a real-world problem [1] that is not based on boson or random-circuit sampling. However, this choice does not constrain the quantum problem that can be used for PoQ. By choosing a sufficiently complex ensemble of experiments and sampling it cryptographically, experimental outcomes can be made weakly correlated. Strategies, such as randomizing problem topology, dimensionality, qubit count, and annealing time, can prevent classical learning algorithms from predicting expectation values based on large datasets [50]. Additionally, hash algorithms leveraging entanglement witnesses and shadow tomography can further enhance the quantum nature of the hashing process. We note that while supremacy is desirable for PoQ, it may not be strictly necessary as long as classical competition remains prohibitively expensive, making quantum approaches the only practicable ones. In fact, one may deliberately use smaller problem instances where the ground truth can be obtained — albeit with significant computational effort — to benchmark and calibrate QPUs.
@@
A quantum blockchain could transform existing blockchain technology by significantly mitigating the substantial electricity demands of classical proof of work (PoW) mining. Electricity costs are estimated to comprise 90% to 95% of Bitcoin’s total mining expenses [51]. In our PoQ approach, energy consumption amounts to a small fraction of quantum computation cost (which is the dominant cost in the PoQ approach). For example, energy expenses account for just 0.1% of the total computation cost on D-Wave quantum computers [52]. Replacing classical PoW with quantum PoQ could therefore reduce energy costs by a factor of 1,000. This ratio may shift as quantum computation becomes more cost-efficient over time, but substantial energy savings are expected to persist. Transition to PoQ would also relocate mining from regions with low electricity costs to countries with advanced quantum computing infrastructure. Quantum computation has the potential to undermine the security of some proof-of-work blockchains, unstructured database search (Grover’s algorithm) being the best known example [53]. Unitary evolution in the quantum-supremacy regime is at the core of our algorithm; Grover’s algorithm requires an efficient oracle to determine membership, which cannot be constructed in our case, and thus a layer of security is provided. Beyond demonstrating the first practical use of quantum supremacy in a critical domain, this work proves the feasibility of quantum computing in real-world tasks and suggests broader high-impact applications with today’s technology.
@@ Appendix A Cross-Validation Limitations in Boson and Random-Circuit Sampling
To operate within a quantum blockchain, QPUs must cross-validate each other’s outcomes [27]. Unstructured distributions yielding states with close to uniform random probability, such as those obtained from random-circuit sampling, are unsuitable for this purpose: the expectation values of all foreseeable k-local operators remain close to zero, making them ineffective in the construction of witnesses. The only observable with a nontrivial expectation is $\left\langle\psi\right|\rho\left|\psi\right\rangle$, where $\left|\psi\right\rangle$ is a ground-truth wave function. For this reason the estimator commonly used to establish supremacy is the cross entropy [38]. Given a distribution allowing fair and efficient sampling in the computational basis ($P_{\text{QPU}}$, typically a QPU), and a well-resolved ground-truth distribution in the same basis ($P_{\text{GT}}$, used for verification), the cross entropy is quantified as
@@
$P_{\text{GT}}$ is typically computed using tensor-network techniques at small scales. This allows the sum to be evaluated by Monte Carlo sampling (of $P_{\text{QPU}}$, i.e. use of a sample set). Calculating $P_{\text{GT}}$ by classical methods becomes intractable at large system sizes (thence supremacy). The alternative is then to determine $P_{\text{GT}}$ from a QPU. XEB is small in experimental demonstrations, even using an ideal (tensor network) distribution as $P_{GT}$ [32, 33]. Nevertheless, we might hope that $P_{GT}$ could be constructed for the purpose of validating quantum work.
@@
A.1 Coarse-grained boson sampling
A proposal for a proof-of-work blockchain using coarse-grained boson sampling (CGBS) was recently proposed by Singh et al. [29]. The form of unitary evolution and measurement that is undertaken is defined by boson sampling. In order to bypass the problem of resolving state probabilities, samples are binned; the statistic to be cross validated is then the most-probable bin, which with care and sufficiently (polynomially many) samples may be reproducible given a set of devices of comparable (high) control [27, 54]. Miners’ broadcasts are synchronized and staked (a miner pays per broadcast), and results are collected subject to a classical-filtering process (mode binning) into a single ground truth. If miners submit sample sets well-matched to the ground truth they are rewarded in proportion to the number of samples broadcast (work). Useful bounds are derived as a function of sampling error and control (photon and other losses), making clear that further development of these devices is necessary for an experimental implementation.
@@ A.1 Coarse-grained boson sampling
The consensus mechanism based on CGBS faces several challenges including an absence of devices operating with quantum supremacy and the need to improve quality of device. The collective ground-truth consensus mechanism proposed is quite different from the Bitcoin-like scheme we demonstrate. Whereas the evaluation of work in our proposal requires that a stakeholder trust their own verifying device, in the CGBS proposal there is a need to trust an aggregated ground truth across many miners built with a requirement for strong network synchronization. Several interesting mechanisms are proposed to mitigate for weakness of the protocol, some similar to those in Section C. In particular we note that with the combination of a stake and an efficient classical filter, many obvious attacks, based on manipulation of the ground truth, become impracticable. Further development may be require to mitigate for attacks that involve recombining, resubmitting and/or buffering samplesets for enhanced reward (attacks which even when unrewarded may destabilize verification). Note that states might be manipulated in a manner preserving exact or approximate mode populations. There are also weaknesses in our blockchain arising from, for example, the redefinition of Chainwork that do not impact this proposal.
@@ B.1 Multi-Basis Measurement
In some quantum systems, it is possible to perform measurements in multiple bases. This is feasible, for example, when the system allows single-qubit rotations (whether Clifford or non-Clifford gates) prior to measurement. Both digital and digital-analog QPUs can potentially enable such capabilities. Measuring in multiple bases allows access to off-diagonal elements of the density matrix, $\rho_{\mathcal{M}}$. Given their exponential growth in number, accessing all off-diagonal elements is both impractical and unnecessary for most applications. Therefore, rather than attempting to recover all elements of the density matrix, one can measure quantities accessible through sampling by performing measurements in multiple bases.
@@ B.2 CHSH Inequality
[…]
@@
Since the number of pairs is $N(N-1)/2$, the equality is satisfied if $\langle\text{CHSH}\rangle_{\rho_{ij}}=2$ for all pairs. This means that it is not possible to violate all CHSH inequalities at the same time. In other words, if some of the inequalities are violated, there are always others that are not violated. This is closely related to the monogamy of entanglement, which states that if two subsystems are strongly entangled, they should have less entanglement with other subsystems.
@@ Appendix C Weaknesses and Mitigations
As described in the main text, the particular proof-of-concept experiment is premised on the notion of honest network users and is susceptible to attacks. In addition to mitigations described in the main text we explore additional procedures in this appendix. We present a set of tools to protect against no-work and partial-work attacks, but optimization of the experimental ensemble, post-processing and digitalization process, essential in practice, goes beyond the scope of this publication.
C.1 No Work Attacks
We consider first the possibility that a miner might present a fraudulent block, one in which the nonce is chosen uniformly at randomly without conducting any experiment. In Bitcoin a random nonce can be used in a broadcast block. This convinces either all users with probability $2^{-\mathcal{N}_{\rm zeros}}$ or convinces no users. The expected reward from such a strategy is $2^{-\mathcal{N}_{\rm zeros}}R-C_{B}-C_{P}$, where $C_{B}$ is the cost of broadcasting the block, $R$ is the expected block reward and $C_{P}$ the cost of block preparation. Pursuing this strategy repeatedly the cost of block preparation is negligible (only a bit increment of the nonce is required between proposals). By contrast the expected reward from honest mining is $((R-C_{B})2^{-\mathcal{N}_{\rm zeros}}-C_{P}-C_{H})$, where $C_{H}$ is the cost of hashing (validation). The reward in Bitcoin is large enough compared to the broadcast cost and $C_{H}$ that mining is incentivized. Simultaneously $C_{B}>2^{-\mathcal{N}_{\rm zeros}}R$, so random broadcasts are disincentivized.
@@ C.1 No Work Attacks
DoS and random nonce mitigations effective in other mature blockchain technologies can be used (see e.g.  [59]). The situation is qualitatively the same in our probabilistic blockchain, although the expected reward is smaller as a function of the blockchain efficiency (enhanced branching), and validation may be inhomogeneous with respect to users. Validation is also more expensive than in many other blockchains, so fewer malicious packets can be more disruptive. The inhomogeneity of validation is a problem, since mining power may be diluted by such an attack even if the attacker is not ultimately rewarded.
@@
C.2 Partial work attacks
@@
Partial work does not offer attack opportunities in Bitcoin owing to the cryptographic properties of SHA-256. A hypothetical probabilistic hash might fail in an independent and identically distributed manner, providing similar guarantees against partial work. However, our hash is to be inferred from a known experimental protocol ($\Theta$) and digitalization ($G,W_{0}$) procedure, which in practical implementations can leave room for inference to bypass at least some quantum work.
@@
Consider the case of a blockchain built upon quantum supremacy in the estimation of quench statistics [1], used in our proof of concept. Since signs on hyperplanes are determined as a (classical) cryptographic function of the block header, preimage resistance is assured: it is not possible to determine a block header from a desired (all zero) bit pattern. However, given a nonce (thence protocol parameters $\Theta,G,W_{0}$) one can estimate correlations with increasing accuracy by a hierarchy of classical computational methods (of increasing cost). Certain experiments ($\Theta$) may allow better or worse estimators, combined with certain realizations of $G$ and $W_{0}$ that allow an accumulation of signal relative to the error. In this way the sign $W_{\alpha}-W_{0}$ for some (many) bits might be inferred with non-negligible (or perhaps even high) confidence.
@@
An attack might proceed as follows. For many nonces an attacker can estimate $\{W_{\alpha}\}$ by classical methods (e.g. by approximate simulation, or machine learning). The results are ranked by the number of (robust) threshold-satisfying bits. The best candidate might be broadcast directly. The reward from this strategy is described as in the no-work scenario, but with an enhanced (classical) work cost and enhanced probability that the bits are verified. This attack is mitigated in part by the modified block structure of Appendix C.4. In particular, spoofing enhanced experimental data (beyond the $\mathcal{N}_{\rm zeros}$ proof of work bits) is impossible given quantum supremacy [1].
@@
A second way to use classical estimators is as a filter. Rather than broadcast the most promising nonce, the nonces are ranked and subjected to the prescribed quantum experiment. Only if the quantum work validates the nonce is it broadcast. The filter allows a user to bypass experiments that are less likey to meet the Chainwork requirement. In this case, the filter success rate must grow slowly enough with the classical work for quantum experiments to be favored over the filter.
@@
To mitigate this filter attack it is necessary that all practical estimators are weak. One simple mechanism by which to weaken classical estimators is described in Appendix C.5. The quality of estimators as a function of the classical-work investigated is above all a function of the experimental ensemble (distribution of $\Theta$).

### v3

@@ VI Demonstrated Implementation
Fig. 5 illustrates the mining efficiency across multiple blockchain operations. Mining efficiency is defined as the fraction of block broadcasts, excluding the genesis block, that successfully enter the strongest chain in the limit of many miners. For each curve the data is taken from 16 chains created with bootstrapping of witness distributions, each having between 512 and 1024 block broadcasts. The bootstrapping method is in agreement with results without bootstrapping reported in other figures. In Appendix D.1 we describe mining efficiency and transaction delay in greater detail, presenting additional data. The analysis compares different Chainwork definitions and post-processing methods. For the basic $\pm 1$ weighting method, which assigns equal weights to all block broadcasts, mining efficiency drops to close to 50% as $\mathcal{N}_{\rm zeros}\to 64$. The confidence-based weighting technique performs better, as expected, approaching 75% efficiency for $\mathcal{N}_{\rm zeros}=64$. Confidence-based Chainwork allows higher efficiency than basic +/-1 Chainwork, also when accounting for small differences in the mining rate as outlined in Appendix D.2. To calculate confidence values we used a constant $\delta W_{\alpha}$ extracted from demonstration data, as detailed in Appendix D.
We would emphasize that per nonce there is a unitary evolution (QPU programming), which in our proof of concept requires 1 second of QPU access time on a 12.5kW quantum computer [53]. Per nonce there is also a need to perform one classical cryptographic hash for generation of unitary parameters, amongst other low-cost classical computations (generation of unitary evolution parameters, summation of read-out states for statistical estimation, and locality-sensitive hashing [projection of statistics onto witnesses by random vectors]). If the cryptographic hash is calculated on a modern CPU, for example SHA-256, this requires of the order of several hundred clock cycles, or approximately 100 ns[54]. Even if one were to consider shorter (but still practically possible) quantum demonstrations, the classical costs per nonce attempt (including the classical hash) are negligible by comparison with operation of the quantum computer in terms of time, energy and dollars.
In this context we can also highlight significant advantages in terms of energy consumption by comparison with classical proof-of-work blockchains operating at equivalent block value. A quantum computer can cost on the order of 10 million US dollars and operate on the order of years[55]. The cost is therefore of the order of thousands of USD per hour, by contrast with the energy cost of 12.5kW [53], which is on the order of 1 USD per hour. The dollar value of computer power brought to bear in mining is expected to scale with the value of the block in both classical and quantum blockchains, if sufficient supply can be achieved. Electricity costs are estimated to comprise 90% to 95% of Bitcoin’s total mining expenses [51]. In quantum computing the relative energy cost would be of the order of $O(1/1000)$.
In combination with a reference Python implementation, witness distributions for the four QPUs used in this study for all resampling analyses are available in Zenodo at [52]. The reference implementation allows construction of blockchains with either generally-available QPUs, or reproduction of the bootstrapping methodology using the data collected for this study.
@@ VII Conclusions
We have proposed and successfully tested a blockchain architecture in which proof of quantum work (PoQ) is performed on quantum computers. Our approach introduces multiple quantum-based hash generation methods, with varying levels of complexity depending on the capabilities of the quantum hardware. To address the inherent probabilistic nature of quantum hash generation and validation, an unavoidable property of any quantum hash algorithm, we developed techniques to ensure blockchain stability. We tested our proposed blockchain on four D-Wave annealing QPUs, from two processor generations located in the United States and Canada. We demonstrated that despite architectural and parametric variations, these QPUs can be programmed to generate and validate each other’s hashes in a way that allowed stable blockchains to persist over thousands of broadcast blocks and hundreds of thousands of hash validations.
@@
For this demonstration, we used the critical dynamics of 3D and biclique spin glass at a fixed size as the computational problem for PoQ. This was motivated by its role in the first demonstration of beyond-classical computation in a real-world problem [1] that is not based on boson or random-circuit sampling. However, this choice does not constrain the quantum problem that can be used for PoQ. By choosing a sufficiently complex ensemble of unitary evolutions, in combination with sufficient sampling and post-processing, statistical estimation to high precision by classical computers is made infeasible. Strategies, such as randomizing problem topology, dimensionality, qubit count, and annealing time, can prevent classical learning algorithms from predicting expectation values based on large datasets [50]. Additionally, hash algorithms leveraging entanglement witnesses and shadow tomography can further enhance the quantum nature of the hashing process. We note that while beyond-classical computation is desirable for PoQ, it may not be strictly necessary as long as classical competition remains prohibitively expensive, making quantum approaches the only practicable ones. In fact, one may deliberately use smaller problem instances where the ground truth can be obtained — albeit with significant computational effort — to benchmark and calibrate QPUs.
@@
A quantum blockchain could transform existing blockchain technology by significantly mitigating the substantial electricity demands of classical proof-of-work (PoW) mining. In our PoQ approach, energy consumption amounts to a small fraction of quantum computation cost (which is the dominant cost in the PoQ approach). For example, energy expenses account for just 0.1% of the total computation cost on D-Wave quantum computers [53]. Replacing classical PoW with quantum PoQ could therefore reduce energy costs by a factor of 1,000. This ratio may shift as quantum computation becomes more cost-efficient over time, but substantial energy savings are expected to persist. Transition to PoQ would also relocate mining from regions with low electricity costs to countries with advanced quantum computing infrastructure. Quantum computation has the potential to undermine the security of some proof-of-work blockchains, unstructured database search (Grover’s algorithm) being the best-known example [56]. Unitary evolution in the beyond-classical regime is at the core of our algorithm; Grover’s algorithm requires an efficient oracle to determine membership, which cannot be constructed in our case, and thus a layer of security is provided. Beyond demonstrating the first practical use of beyond-classical computation in a critical domain, this work proves the feasibility of quantum computing in real-world tasks and suggests broader high-impact applications with today’s technology.
Data availability
A limited set of experimental data is available in the Zenodo online repository, as well as a reference software implementation of methods[52].
@@ Appendix A Cross-Validation Limitations in Boson and Random-Circuit Sampling
To operate within a quantum blockchain, QPUs must cross-validate each other’s outcomes [27]. Unstructured distributions yielding states with close to uniform random probability, such as those obtained from random-circuit sampling, are unsuitable for this purpose: the expectation values of all foreseeable k-local operators remain close to zero, making them ineffective in the construction of witnesses. The only observable with a nontrivial expectation is $\left\langle\psi\right|\rho\left|\psi\right\rangle$, where $\left|\psi\right\rangle$ is a ground-truth wave function. For this reason the estimator commonly used to establish beyond-classical computation is the cross entropy [38]. Given a distribution allowing fair and efficient sampling in the computational basis ($P_{\text{QPU}}$, typically a QPU), and a well-resolved ground-truth distribution in the same basis ($P_{\text{GT}}$, used for verification), the cross entropy is quantified as
@@
$P_{\text{GT}}$ is typically computed using tensor-network techniques at small scales. This allows the sum to be evaluated by Monte Carlo sampling (of $P_{\text{QPU}}$, i.e. use of a sample set). Calculating $P_{\text{GT}}$ by classical methods becomes intractable at large system sizes. The alternative is then to determine $P_{\text{GT}}$ from a QPU. XEB is small in experimental demonstrations, even using an ideal (tensor network) distribution as $P_{GT}$ [32, 33]. Nevertheless, we might hope that $P_{GT}$ could be constructed for the purpose of validating quantum work.
@@
A.1 Coarse-grained boson sampling blockchains
A proposal for a proof-of-work blockchain using coarse-grained boson sampling (CGBS) was recently proposed by Singh et al. [29]. The form of unitary evolution and measurement that is undertaken is defined by boson sampling. In order to bypass the problem of resolving state probabilities, samples are binned; the statistic to be cross validated is then the most-probable bin, which with care and sufficiently (polynomially many) samples may be reproducible given a set of devices of comparable (high) control [27, 57]. Miners’ broadcasts are synchronized and staked (a miner pays per broadcast), and results are collected subject to a classical-filtering process (mode binning) into a single ground truth. If miners submit sample sets well-matched to the ground truth they are rewarded in proportion to the number of samples broadcast (work). Useful bounds are derived as a function of sampling error and control (photon and other losses), making clear that further development of these devices is necessary for a demonstration.
@@ A.1 Coarse-grained boson sampling blockchains
The consensus mechanism based on CGBS faces several challenges including an absence of devices operating in the beyond-classical regime, and the need to improve quality of device. The collective ground-truth consensus mechanism proposed is quite different from the Bitcoin-like scheme we demonstrate. Whereas the evaluation of work in our proposal requires that a stakeholder trust their own verifying device, in the CGBS proposal there is a need to trust an aggregated ground truth across many miners built with a requirement for strong network synchronization. Several interesting mechanisms are proposed to mitigate for weakness of the protocol, some similar to those in Section C. In particular we note that with the combination of a stake and an efficient classical filter, many obvious attacks, based on manipulation of the ground truth, become impracticable. Further development may be require to mitigate for attacks that involve recombining, resubmitting and/or buffering samplesets for enhanced reward (attacks which even when unrewarded may destabilize verification). Note that states might be manipulated in a manner preserving exact or approximate mode populations. There are also weaknesses in our blockchain arising from, for example, the redefinition of Chainwork that do not impact this proposal.
@@ B.1 Multi-Basis Measurement
In some quantum systems, it is possible to perform measurements in multiple bases. This is feasible, for example, when the system allows single-qubit rotations (whether Clifford or non-Clifford gates) prior to measurement. Both digital and digital-analog QPUs can potentially enable such capabilities. Measuring in multiple bases allows access to off-diagonal elements of the density matrix, $\rho_{\mathcal{M}}$. Given their exponential growth in number, accessing all off-diagonal elements is both impractical and unnecessary for most applications. Therefore, rather than attempting to recover all elements of the density matrix, one can measure quantities accessible through sampling by performing measurements in multiple bases. Digital-analog methods can be realized in experimental D-Wave quantum annealing processors using multi-color annealing methods for purposes of measuring non-computational basis information[58].
@@ B.2 CHSH Inequality
where $\left|\uparrow\right\rangle$ and $\left|\downarrow\right\rangle$ are eigenstates of $\sigma^{z}$ with eigenvalues $\pm 1$, respectively. The pure state density matrix is
@@
Since the number of pairs is N(N1)/2N(N-1)/2, the equality is satisfied if $\langle\text{CHSH}\rangle_{\rho_{ij}}=2$ for all pairs. This means that it is not possible to violate all CHSH inequalities at the same time. In other words, if some of the inequalities are violated, there are always others that are not violated. This is closely related to the monogamy of entanglement, which states that if two subsystems are strongly entangled, they should have less entanglement with other subsystems.
@@ Appendix C Weaknesses and Mitigations
As described in the main text, the particular proof-of-concept demonstration is premised on the notion of honest network users and is susceptible to attacks. In addition to mitigations described in the main text we explore additional procedures in this appendix. We present a set of tools to protect against no-work and partial-work attacks, but optimization of the ensemble of unitary dynamics, post-processing and digitalization process, essential in practice, goes beyond the scope of this publication.
C.1 No-Work Attacks
We consider first the possibility that a miner might present a fraudulent block, one in which the nonce is chosen uniformly at randomly without conducting any quantum computation or inference. In Bitcoin a random nonce can be used in a broadcast block. This convinces either all users with probability $2^{-\mathcal{N}_{\rm zeros}}$ or convinces no users. The expected reward from such a strategy is $2^{-\mathcal{N}_{\rm zeros}}R-C_{B}-C_{P}$, where $C_{B}$ is the cost of broadcasting the block, $R$ is the expected block reward and $C_{P}$ the cost of block preparation. Pursuing this strategy repeatedly the cost of block preparation is negligible (only a bit increment of the nonce is required between proposals). By contrast the expected reward from honest mining is $((R-C_{B})2^{-\mathcal{N}_{\rm zeros}}-C_{P}-C_{H})$, where $C_{H}$ is the cost of hashing (validation). The reward in Bitcoin is large enough compared to the broadcast cost and $C_{H}$ that mining is incentivized. Simultaneously $C_{B}>2^{-\mathcal{N}_{\rm zeros}}R$, so random broadcasts are disincentivized.
@@ C.1 No-Work Attacks
DoS and random nonce mitigations effective in other mature blockchain technologies can be used (see e.g.  [63]). The situation is qualitatively the same in our probabilistic blockchain, although the expected reward is smaller as a function of the blockchain efficiency (enhanced branching), and validation may be inhomogeneous with respect to users. Validation is also more expensive in our simplest (main-text) implementation, so fewer malicious packets can be more disruptive. The inhomogeneity of validation is a problem, since mining power may be diluted by such an attack even if the attacker is not ultimately rewarded.
@@
We note that if miners are not anonymous their behavior can be tracked. A rule can be introduced to penalize or temporarily block users who consistently propose unverifiable (or statistically anomalous) work, as is done for DoS. Note that a transmission path or node might be penalized with limited information on the stakeholder. We will later see that deanonymization is not essential to mitigate work and partial work attacks.
C.2 Partial-work attacks
@@
Partial work does not offer attack opportunities in Bitcoin owing to the cryptographic properties of SHA-256. A hypothetical probabilistic hash might fail in an independent and identically distributed manner, providing similar guarantees against partial work. However, our hash is to be inferred from a known unitary evolution ($\Theta$) and digitalization ($G,W_{0}$) procedure, which in practical implementations can leave room for inference to bypass at least some quantum work.
@@
Consider the case of a blockchain built upon advantage in the estimation of quench statistics [1], used in our proof of concept. Since signs on hyperplanes are determined as a (classical) cryptographic function of the block header, preimage resistance is assured: it is not possible to determine a block header from a desired (all zero) bit pattern. However, given a nonce (thence protocol parameters $\Theta,G,W_{0}$) one can estimate correlations with increasing accuracy by a hierarchy of classical computational methods (of increasing cost). Certain unitary evolutions ($\Theta$) may allow better or worse estimators, combined with certain realizations of $G$ and $W_{0}$ that allow an accumulation of signal relative to the error. In this way the sign $W_{\alpha}-W_{0}$ for some (many) bits might be inferred with non-negligible (or perhaps even high) confidence.
@@
An attack might proceed as follows. For many nonces an attacker can estimate $\{W_{\alpha}\}$ by classical methods (e.g. by approximate simulation, or machine learning). The results are ranked by the number of (robust) threshold-satisfying bits. The best candidate might be broadcast directly. The reward from this strategy is described as in the no-work scenario, but with an enhanced (classical) work cost and enhanced probability that the bits are verified. This attack is mitigated by the modified block structure of Appendix C.4. In particular, spoofing enhanced experimental data is impossible given beyond-classical computation [1]; we demonstrate methods to cryptographically secure and decouple the computational complexity requirement $\mathcal{N}_{\rm zeros}$ from the length of the quantum hash allowing comparison of higher precision experimental results.
@@
A second way to use classical estimators is as a filter. Rather than broadcast the most promising nonce, the nonces are ranked and subjected to the prescribed quantum experiment. Only if the quantum work validates the nonce is it broadcast. The filter allows a user to bypass experiments that are less likely to meet the Chainwork requirement. In this case, the filter success rate must grow slowly enough with the classical work for quantum experiments to be favored over the filter.
@@
A modified block structure can also mitigate for the filter attack as later described. To mitigate this filter attack without a modified block structure it is necessary that all practical estimators are weak. One simple mechanism by which to weaken classical estimators is described in Appendix C.5. The quality of estimators as a function of the classical-work investigated is above all a function of the unitary-evolution ensemble (distribution of $\Theta$); it is important to choose a unitary evolution with a maximum gap between classical and quantum estimators.